Given the μ=65 and σ=5 of a population of Math Exam scores. The z−value that corresponds to a score of X=74 is



1
Expert's answer
2022-02-15T17:32:41-0500

Solution:

 μ=65 and σ=5

X=74

"z=\dfrac{x-\mu}{\sigma}
\\\Rightarrow z=\dfrac{74-65}{5}=1.8"